OpenStudy (anonymous):

-8x - 8y = -16, 6x - 9y = -108

OpenStudy (tkhunny):

Make your life easier! -8x -8y = -16 ==> -2x - 2y = -4 6x - 9y = -108 ==> 2x - 3y = -36